OpenStudy (anonymous):

Write 18/12 in lowest terms. Express your answer as an improper fraction.

OpenStudy (triciaal):

FACTORS OF 18/FACTORS OF 12

OpenStudy (anonymous):

You need to start finding numbers that can be divided my both numbers. 2 is always a good guess 18/2 = 9 12/2 = 6 So 18/12 = (18/2)/(12/2) = 9/6 Now you need to find a number that ca be divided by both 9 and 6... 3 is a good guess. 9/3 = 3 6/3 = 2 9/6 = (9/3)/(6/3) = 3/2 Now you cant simply continue because there is any other number capable of being divided by both 2 and 3 (apart from 1) so this is the lowest term possible. Hope it helps

OpenStudy (anonymous):

Thank you so much